OpenStudy (anonymous):

What does Descartes's Rule of Signs tell you about the real roots of the polynomial? –2x^3 + 3x^2 – 5x – 2 = 0 There is one positive root and either 2 or 0 negative roots. There are either 2 or 0 positive roots and there are either 2 or 0 negative roots. There is one positive root and one negative root. There are either 2 or 0 positive roots and one negative root.

OpenStudy (anonymous):

(2x + 1)(x + 1) = 2x^2 + 3x + 1 (2x^2 + 3x + 1)(x - 2) = 2x^3 -x^2 - 5x - 2) (2x + 1)(x + 1)(x - 2) = 0 x = 2, -1, -1/2

OpenStudy (anonymous):

so is it c ?

OpenStudy (anonymous):

Yeah its c

OpenStudy (anonymous):

thanks (; can you help me w/this question here : http://openstudy.com/users/andreascott#/updates/50906154e4b0ad6205378bbf

OpenStudy (anonymous):

yeah I'll try